# On the Sylvester program and Cayley algorithm for vector partition reduction

### v2

@@ 3.3 General Cayley method for P- and NP-columns
The original Cayley reduction method presented in Section 3.1 applied to double partition with the matrix ${\bf D}$ having $m$ columns generates a sum (13) of $m$ SPFs. This approach fails when at least one of the columns, say the first column ${\bf c}_{1}$, has GCD of its elements larger than unity $\gcd({\bf c}_{1})=\gcd(b_{1},\beta_{1})=g_{1}>1,$ and we write $b_{1}=g_{1}b^{\star},\ \beta_{1}=g_{1}\beta^{\star}$. This case should be treated as discussed above in Section 3.2 leading to $C_{r,\rho}^{1}$ in (24).
@@
Consider (24) and note that when $\gcd(b_{1},\beta_{1})=g_{1}=1$ the symbolic argument in SPFs in the summand is an integer only if the condition $j_{y}=\rho\bmod{\beta_{1}}$ holds. This means that when ${\bf c}_{1}$ is a P-column the condition $j_{y}=\rho\bmod{\beta_{1}}$ might be dropped and all terms in the sum in (24) contribute to $C_{r,\rho}^{1}$ that has an alternative compact form used in (13).
@@
In case of NP-column ${\bf c}_{1}$ with $g_{1}>1$ we have $(\rho-j_{y})b_{1}/\beta_{1}=(\rho-j_{y})b^{\star}/\beta^{\star},$ and for given value of $\rho$ there exist exactly $g_{1}$ values of $j_{y}$ for which the SPF symbolic argument $r-j_{x}-(\rho-j_{y})b_{1}/\beta_{1}$ has integer value. The condition $j_{y}=\rho\bmod{\beta_{1}}$ selects only one of $g_{1}$ values of $j_{y}$ that should be used for the estimate of the contribution to VPF. Thus not all (but each $g_{1}$-th only) summands with integer SPF symbolic argument contribute and this fact prevents the compactification of the superposition (24) into a single term. An example of NP-column reduction in double partition is discussed in Appendix A.
3.4 Zero in last row of matrix ${\bf D}$
@@
Note that for $\beta_{1}=0$ the expression for $S_{1}(x,y_{0})$ in (18) vanishes implying vanishing contribution $C^{1}_{r,\rho}=0$. We also observe that the result (15) with $i=1$ and $\beta_{1}=0$ formally produces all $d_{1j}<0$ and $L_{1}=-s_{2}b_{1}$ leading to negative $L_{1}({\bf s})=-s_{2}b_{1}+\sum_{j=2}^{m}d_{1j}$ for any positive $s_{2}$. This means that the scalar partition $W(L_{1},{\bf d}_{1})=W(L_{1}({\bf s}),|{\bf d}_{1}|)$ equals zero. Thus the rule of thumb for double partitions with nonnegative generators $b_{i},\beta_{i}$ and $r,\rho\geq 0$ is – the contribution of the column with $\beta_{i}=0$ is zero.
There is another possibility to evaluate a double partition with ${\bf c}_{1}=\{b_{1},0\}^{T}$. Rewrite the system ${\bf s}={\bf D\cdot x}$ as follows
${\bf s}={\bf c}_{1}x_{1}+\sum_{i=2}^{m}{\bf c}_{i}x_{i}\Rightarrow{\bf s}^{\prime}_{1}=\sum_{i=2}^{m}{\bf c}_{i}x_{i},\quad{\bf s}^{\prime}_{1}={\bf s}-{\bf c}_{1}x_{1}=\{r-b_{1}x_{1},\rho\}^{T}.$ (27)
Introduce $\bm{x}^{\prime}_{1}=\{x_{2},x_{3},\ldots,x_{m}\}$ and use the vectors $\bm{b}^{\prime}_{1},\ \bm{\beta}^{\prime}_{1}$ in (19) to obtain a system of two equations
$\bm{b}^{\prime}_{1}\cdot\bm{x}^{\prime}_{1}=r-b_{1}x_{1},\quad\bm{\beta}^{\prime}_{1}\cdot\bm{x}^{\prime}_{1}=\rho.$
The matrix $\bm{D}^{\prime}_{1}=\{{\bf c}_{2},\ldots,{\bf c}_{m}\}=\{\bm{b}^{\prime}_{1},\bm{\beta}^{\prime}_{1}\}$ made of the rows $\bm{b}^{\prime}_{1}$ and $\bm{\beta}^{\prime}_{1}$ together with the column ${\bf s}^{\prime}_{1}$ define a double partition $W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})$ for each nonnegative value $x_{1}\geq 0$. This means that
$W({\bf s},{\bf D})=\sum_{x_{1}=0}^{\infty}W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1}).$
The relations (13) determine the contribution $W_{i}^{2}({\bf s})$ obtained by elimination of the P-column ${\bf c}_{i},\ i>1$ as $W(L_{i},{\bf d}_{i})$ where ${\bf d}_{i}=\{d_{i1},d_{i2},d_{i3},\ldots,d_{im}\}$ and $d_{i1}=b_{1}\beta_{i}$. The same reduction procedure for $W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})$ leads to the term $W(L_{i}-x_{1}b_{1}\beta_{i},\bm{d}^{\prime}_{i})=W(L_{i}-x_{1}d_{i1},\bm{d}^{\prime}_{i})$ with $\bm{d}^{\prime}_{i}=\{d_{i2},d_{i3},\ldots,d_{im}\}$. The definition (2) of the generating function for $W(s,{\bf d})$ implies the recursion
$W(s,{\bf d})-W(s-d_{1},{\bf d})=W(s,\{d_{2},d_{3},\ldots,d_{m}\}).$ (28)
Estimate the infinite sum of the terms $W(L_{i}-x_{1}d_{i1},\bm{d}^{\prime}_{i})$ with $i\geq 0$ to obtain
$\sum_{x_{1}=0}^{\infty}W(L_{i}-x_{1}d_{i1},\bm{d}^{\prime}_{i})=\sum_{x_{1}=0}^{\infty}[W(L_{i}-x_{1}d_{i1},{\bf d}_{i})-W(L_{i}-(x_{1}+1)d_{i1},{\bf d}_{i})]=W(L_{i},{\bf d}_{i}),$ (29)
as there always exists a positive integer $k$ such that $W(L_{i}-kd_{i1},{\bf d}_{i})$ evaluates to zero. Now write the double partition $W({\bf s},{\bf D})$ as
$W({\bf s},{\bf D})=C_{r,\rho}^{1}+\sum_{i=2}^{m}W(L_{i},{\bf d}_{i}),$
where $C_{r,\rho}^{1}$ denotes (yet unknown) contribution of the column ${\bf c}_{1}$. The relation (29) shows that each contribution $W(L_{i},{\bf d}_{i})$ for $2\leq i\leq m$ can be written as the infinite sum of $W(L_{i}-x_{1}d_{i1},\bm{d}^{\prime}_{i})$. This means that
$W({\bf s},{\bf D})=\sum_{x_{1}=0}^{\infty}W({\bf s}^{\prime}_{1},\bm{D}^{\prime}_{1})=\sum_{i=2}^{m}\sum_{x_{1}=0}^{\infty}W(L_{i}-x_{1}d_{i1},\bm{d}^{\prime}_{i})=\sum_{i=2}^{m}W(L_{i},{\bf d}_{i}),$ (30)
and the contribution $C_{r,\rho}^{1}$ indeed equals zero. This property is used in an example discussed in Appendix B.
Note that (29) is valid only for $b_{1}>0$, nevertheless the case $b_{1}<0$ can be reduced to the problem with positive $b_{1}$. Rewrite the generating function (8) by multiplying both the numerator and denominator by $x^{-b_{1}}$ and obtain
$G({\bf t},{\bf D})=\frac{x^{-b_{1}}}{(1-x^{-b_{1}})}\prod_{i=2}^{m}(1-x^{b_{i}}y^{\beta_{i}})^{-1}.$
This expression corresponds to the double partition $W({\bf s}^{\prime},{\bf D}^{\prime})$ with
${\bf D}^{\prime}=\{-{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\},\quad{\bf s}^{\prime}={\bf s}+{\bf c}_{1}=\{r+b_{1},\rho\}^{T},$
we have $W({\bf s},{\bf D})=W({\bf s}^{\prime},{\bf D}^{\prime})$ and the contribution of the first column of ${\bf D}^{\prime}$ to $W({\bf s}^{\prime},{\bf D}^{\prime})$ is zero.
The matrix $\bm{D}^{\prime}_{1}$ used in derivation of (30) is assumed to be free of zeros in its last row. When this matrix has another column ${\bf c}_{j}$ with second zero element the columns ${\bf c}_{1}$ and ${\bf c}_{j}$ are parallel to each other and we refer the reader to Section 5 that considers the general case of collinear columns.
@@ 4 $l$-tuple partition reduction
Consider a partition $W({\bf s},{\bf D})$ for a non-negative $l\times m\ (l<m)$ matrix ${\bf D}=\{{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of $m$ columns ${\bf c}_{i}=\{c_{i1},c_{i2},\ldots,c_{il}\}^{T}$ in assumption of column noncollinearity. We would also use equivalent notation $W({\bf E})\equiv W({\bf s},{\bf D})$ for the partition with corresponding augmented matrix ${\bf E}=\{{\bf c}_{0}={\bf s},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ obtained by adding the symbolic argument column ${\bf s}$ to matrix ${\bf D}$.
@@ 4.1 Partial fraction expansion
[…]
$S_{1}(\hat{\bf t},t_{l0})=\prod_{i\neq 1}^{m}\prod_{j=1}^{c_{1l}-1}(1-\omega_{1j}\hat{\bf t}^{\hat{\bf c}_{i}}t_{l0}^{c_{il}}).$
The denominator reads
$\Pi_{1}(\hat{\bf t})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}}t_{l0}^{c_{1l}c_{il}})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}-c_{il}\hat{\bf c}_{1}}),$ (34)
while for the numerator we have:
$S_{1}(\hat{\bf t},t_{l0})=\Pi_{1}(\hat{\bf t})A_{1}(\hat{\bf t},t_{l0})=\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l0}^{j_{l}},$ (35)
where $A_{1,j_{l}}$ are rational functions in $\hat{\bf t}$. To find these functions use
$S_{1}(\hat{\bf t},t_{l0})=\prod_{i\neq 1}^{m}\frac{(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}}t_{l,0}^{c_{1l}c_{il}})}{(1-\hat{\bf t}^{\hat{\bf c}_{i}}t_{l0}^{c_{il}})}=\prod_{i\neq 1}^{m}\left(\sum_{k_{i}=0}^{c_{1l}-1}\hat{\bf t}^{k_{i}\hat{\bf c}_{i}}t_{l0}^{k_{i}c_{il}}\right).$ (36)
@@ 4.2.1 Compact form
Recall that $t_{l0}=\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}}$ and using (33,34) find
@@
It should be reminded here that the final goal of the Sylvester program is the VPF reduction to a set of SPFs and thus the elimination of a single (last) row of $l\times m$ matrix is just the first step of the procedure. We discuss the computation aspects of the Sylvester program further in Section 6.4 but now we have to underline that the reduction procedure described above is valid only when the matrix under reduction has last row void of nonpositive elements. It follows from (50) that some columns ${\bf c}^{1}_{j}$ might have negative elements $c^{1}_{jl}<0$ in the last position.
@@ 4.3 Zero in last row of matrix ${\bf D}$
as there always exists a positive integer $k$ such that $W({\bf s}^{i}-k{\bf c}^{i}_{1},{\bf D}_{i})$ evaluates to zero. It is true if at least one element of ${\bf c}^{i}_{1}=c_{il}\hat{\bf c}_{1}$ is positive which is equivalent to the existence of at least one positive element in the column ${\bf c}_{1}$. If this condition fails, we apply the column sign inversion procedure (51) to the column ${\bf c}_{1}$ with corresponding update of the column ${\bf s}\Rightarrow{\bf s}-{\bf c}_{1}$.
Now write the vector partition $W({\bf s},{\bf D})$ as
@@
and the contribution $C_{\bf s}^{1}$ vanishes.
@@ 5 Vector partition with collinear columns
The reduction algorithm described in Section 4 can be applied in absence of collinear columns in the generator matrix. In this Section we demonstrate that when this condition fails a vector partition can be reduced to a superposition of scalar partition convolutions.
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that vectors corresponding to the first $n$ columns of the incompetent generator matrix ${\bf D}$ are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows
@@
where $u_{j}$ are positive integers and vector ${\bf C}$ represents the GCD vector of the collinear vectors ${\bf c}_{i},1\leq i\leq n<m$. The problem (66) is equivalent to
${\bf S}(k)={\bf s}-k{\bf C}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k=\sum_{i=1}^{n}u_{i}x_{i},\quad 0\leq k\leq K=\min(s_{r}/C_{r}),\quad 1\leq r\leq l,$ (67)
where $K$ is computed using only positive $C_{i}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding double partition $W({\bf S}(k),\bar{\bf D})$ admits a reduction to SPFs either as a sum of $W_{i}^{l}$ or a mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (67) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
$W({\bf s},{\bf D})=\sum_{k=0}^{K}W(k,{\bf u})H({\bf s}-k{\bf C})W({\bf s}-k{\bf C},\bar{\bf D}),\quad K=\min(s_{r}/C_{r}),\quad 1\leq r\leq l,$ (68)
@@
This approach can be extended to the case when the matrix ${\bf D}$ has $r>1$ groups of collinear columns. Each $p$-th group is characterized by its GCD vector ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (67) turns into
@@
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}H({\bf S}({\bf k}))W({\bf S}({\bf k}),\bar{\bf D})\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\ K_{p}=\min(s_{i}/C_{p,i}),\ \sum_{\bf k=0}^{\bf K}\equiv\sum_{k_{1}=0}^{K_{1}}\ldots\sum_{k_{r}=0}^{K_{r}}.$ (70)
@@
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}\sum_{k_{0}=0}^{K_{0}}\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}H({\bf S}({\bf k}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{0}=\max(s_{i}/c_{i}),$ (71)
@@
The extreme case arises for $n=m$, when each column of this matrix belongs to some group of the collinear columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
@@
$W({\bf s},{\bf D})=\sum_{{\bf k}=0}^{{\bf K}}\delta_{{\bf S}({\bf k}^{\prime}),k_{r}{\bf C}_{r}}H({\bf S}({\bf k}^{\prime}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{r}=\max(s_{i}/C_{r,i}).$ (73)

### v3

@@ 3.3 Cayley method for P- and NP-columns
The original Cayley reduction method presented in Section 3.1 applied to double partition with the matrix ${\bf D}$ having $m$ columns generates a sum (13) of $m$ scalar partitions. This approach fails when at least one of the columns, say the first column ${\bf c}_{1}$, has GCD of its elements larger than unity $\gcd({\bf c}_{1})=\gcd(b_{1},\beta_{1})=g_{1}>1,$ and we write $b_{1}=g_{1}b^{\star},\ \beta_{1}=g_{1}\beta^{\star}$. This case should be treated as discussed above in Section 3.2 leading to $C_{r,\rho}^{1}$ in (24).
@@
Consider (24) and note that when $\gcd(b_{1},\beta_{1})=g_{1}=1$ the symbolic argument in scalar partitions in the summand is an integer only if the condition $j_{y}=\rho\bmod{\beta_{1}}$ holds. This means that when ${\bf c}_{1}$ is a P-column the condition $j_{y}=\rho\bmod{\beta_{1}}$ might be dropped and all terms in the sum in (24) contribute to $C_{r,\rho}^{1}$ that has an alternative compact form used in (13).
@@
In case of NP-column ${\bf c}_{1}$ with $g_{1}>1$ we have $(\rho-j_{y})b_{1}/\beta_{1}=(\rho-j_{y})b^{\star}/\beta^{\star},$ and for given value of $\rho$ there exist exactly $g_{1}$ values of $j_{y}$ for which the scalar partition symbolic argument $r-j_{x}-(\rho-j_{y})b_{1}/\beta_{1}$ has integer value. The condition $j_{y}=\rho\bmod{\beta_{1}}$ selects only one of $g_{1}$ values of $j_{y}$ that should be used for the estimate of the contribution to the vector partition. Thus not all (but each $g_{1}$-th only) summands with integer scalar partition symbolic argument contribute and this fact prevents the compactification of the superposition (24) into a single term. An example of NP-column reduction in double partition is discussed in Appendix A.
@@
Note that for $\beta_{1}=0$ the expression for $S_{1}(x,y_{0})$ in (18) vanishes implying vanishing contribution $C^{1}_{r,\rho}=0$. We also observe that the result (15) with $i=1$ and $\beta_{1}=0$ formally produces all $d_{1j}<0$ and $L_{1}({\bf s})=-s_{2}b_{1}$ leading to negative $L^{\prime}_{1}({\bf s})=-s_{2}b_{1}+\sum_{j=2}^{m}d_{1j}$ for any positive $s_{2}$. This means that the scalar partition $W(L_{1}({\bf s}),{\bf d}_{1})=W(L^{\prime}_{1}({\bf s}),|{\bf d}_{1}|)$ equals zero. Thus the rule of thumb for double partitions with nonnegative generators $b_{i},\beta_{i}$ and $r,\rho\geq 0$ is – the contribution of the column with $\beta_{i}=0$ is zero.
The same time we have to recall that in the course of the derivation of (15) we explicitly assumed that $\beta_{1}\neq 0$. Nevertheless it is possible to show that in case $\beta_{1}=0$ the contribution $C^{1}_{r,\rho}$ indeed vanishes. It appears to be a particular case of the general result valid for a $l$-tuple partition ($l\geq 2$) with the zero last element $c_{1l}=0$ in the first column ${\bf c}_{1}$ of the $l\times m$ matrix ${\bf D}$. The reader is referred to Section 4.3 where the details of the derivation are presented.
@@ 4 $l$-tuple partition reduction
Consider a partition $W({\bf s},{\bf D})$ for a non-negative $l\times m\ (l<m)$ matrix ${\bf D}=\{{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ made of $m$ columns ${\bf c}_{i}=\{c_{i1},c_{i2},\ldots,c_{il}\}^{T}$ in assumption of column noncollinearity. We would also use the equivalent notation $W({\bf E})\equiv W({\bf s},{\bf D})$ for the partition with corresponding augmented matrix ${\bf E}=\{{\bf c}_{0}={\bf s},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}$ obtained by adding the symbolic argument column ${\bf s}$ to matrix ${\bf D}$.
@@ 4.1 Partial fraction expansion
Following the approach of Section 3.1 write (29) as
$A_{1}(\hat{\bf t},\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}})=S_{1}(\hat{\bf t},t_{l0})/\Pi_{1}(\hat{\bf t}),\quad\Pi_{1}(\hat{\bf t})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}}t_{l0}^{c_{1l}c_{il}})=\prod_{i\neq 1}^{m}(1-\hat{\bf t}^{c_{1l}\hat{\bf c}_{i}-c_{il}\hat{\bf c}_{1}}).$ (30)
The numerator reads
$S_{1}(\hat{\bf t},t_{l0})=\prod_{i\neq 1}^{m}\left(\sum_{k_{i}=0}^{c_{1l}-1}\hat{\bf t}^{k_{i}\hat{\bf c}_{i}}t_{l0}^{k_{i}c_{il}}\right)=\sum_{j_{l}=0}^{c_{1l}-1}A_{1,j_{l}}(\hat{\bf t})t_{l0}^{j_{l}},$ (31)
@@ 4.2.1 Compact form
Recall that $t_{l0}=\hat{\bf t}^{-\hat{\bf c}_{1}/c_{1l}}$ and using (29,30) find
@@
It should be reminded here that the final goal of the Sylvester program is the vector partition reduction to a set of scalar partitions and thus the elimination of a single (last) row of $l\times m$ matrix is just the first step of the procedure. We discuss the computational aspects of the Sylvester program further in Section 6 but now we have to underline that the reduction procedure described above is valid only when the matrix under reduction has its last row void of nonpositive elements. It follows from (45) that some columns ${\bf c}^{1}_{j}$ might have negative elements $c^{1}_{jl}<0$ in the last position.
@@ 4.3 Zero in last row of matrix ${\bf D}$
Consider evaluation of a $l$-tuple partition with $c_{1l}=0$. Rewrite the system ${\bf s}={\bf D\cdot x}$ as
${\bf s}={\bf c}_{1}x_{1}+\sum_{i=2}^{m}{\bf c}_{i}x_{i}\Rightarrow{\bf s}^{\prime}_{1}=\sum_{i=2}^{m}{\bf c}_{i}x_{i},\quad{\bf s}^{\prime}_{1}={\bf s}-{\bf c}_{1}x_{1}=\{s_{1}-c_{11}x_{1},s_{2}-c_{12}x_{1},\ldots,s_{l}\}^{T}.$ (55)
@@
as there always exists a positive integer $k$ such that $W({\bf s}^{i}-k{\bf c}^{i}_{1},{\bf D}_{i})$ evaluates to zero. It is true if at least one element of ${\bf c}^{i}_{1}=c_{il}\hat{\bf c}_{1}$ is positive which is equivalent to the existence of at least one positive element in the column ${\bf c}_{1}$. If this condition fails, we apply the column sign inversion procedure (46) to the column ${\bf c}_{1}$ with the corresponding update of the column ${\bf s}\Rightarrow{\bf s}-{\bf c}_{1}$. Write the vector partition $W({\bf s},{\bf D})$ as
@@
and the contribution $C_{\bf s}^{1}$ vanishes. This result is used in the example discussed in Appendix B.
@@ 5 Vector partition with collinear columns
The reduction algorithm described in Section 4 cannot be applied in case when in the generator matrix ${\bf D}$ has collinear columns as it leads to appearance of the zero columns in the reduced partition.
Call an integer generator matrix competent if its columns (viewed as vectors) are noncollinear and incompetent otherwise. Assume that the vectors corresponding to the first $n$ columns of the incompetent generator matrix ${\bf D}$ are parallel to each other and rewrite the linear system ${\bf D}\cdot{\bf x}={\bf s}$ as follows
@@
where $u_{j}$ are positive integers and vector ${\bf C}$ represents the GCD vector that divides elementwise all the collinear vectors ${\bf c}_{i},1\leq i\leq n<m$. The problem (62) is equivalent to
${\bf S}(k)={\bf s}-k{\bf C}=\sum_{i=n+1}^{m}{\bf c}_{i}x_{i},\quad k=\sum_{i=1}^{n}u_{i}x_{i},\quad 0\leq k\leq K=\min(\lfloor s_{r}/C_{r}\rfloor),\quad 1\leq r\leq l,$ (63)
where $K$ is computed using only positive $C_{r}>0$. Introduce a vector ${\bf u}=\{u_{1},u_{2},\ldots,u_{n}\}$ and a competent matrix with $(m-n)$ columns $\bar{\bf D}=\{{\bf c}_{n+1},\ldots,{\bf c}_{m}\},$ for which the corresponding partition $W({\bf S}(k),\bar{\bf D})$ admits the reduction either as the sum of $W_{i}^{l}$ or the mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$. The number of solutions of the system (63) for given value of $k$ is equal to a number of solutions of the first equation given by the partition $W({\bf S}(k),\bar{\bf D})$ multiplied by a number of solutions of the second equation, i.e., by a scalar partition $W(k,{\bf u})$. Then the vector partition $W({\bf s},{\bf D})$ is equivalent to a convolution
$W({\bf s},{\bf D})=\sum_{k=0}^{K}W(k,{\bf u})H({\bf s}-k{\bf C})W({\bf s}-k{\bf C},\bar{\bf D}),\quad K=\min(\lfloor s_{r}/C_{r}\rfloor),\quad 1\leq r\leq l,$ (64)
@@
This approach can be extended to the case when the matrix ${\bf D}$ has $r>1$ groups of the collinear columns. Each $p$-th group is characterized by its GCD vector ${\bf C}_{p}$ and the $n_{p}$ integers $u_{p,i}$ making up the vector ${\bf u}_{p}=\{u_{p,1},u_{p,2},\ldots,u_{p,n_{p}}\}$. Introduce the vector ${\bf k}=\{k_{1},k_{2},\ldots,k_{r}\}$ and (63) turns into
@@
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}H({\bf S}({\bf k}))W({\bf S}({\bf k}),\bar{\bf D})\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\ K_{p}=\min(\lfloor s_{i}/C_{p,i}\rfloor),\ \sum_{\bf k=0}^{\bf K}\equiv\sum_{k_{1}=0}^{K_{1}}\ldots\sum_{k_{r}=0}^{K_{r}}.$ (66)
@@
$W({\bf s},{\bf D})=\sum_{\bf k=0}^{\bf K}\sum_{k_{0}=0}^{K_{0}}\delta_{{\bf S}({\bf k}),k_{0}{\bf c}}H({\bf S}({\bf k}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{0}=\max(\lfloor s_{i}/c_{i}\rfloor),$ (67)
@@
The extreme case arises for $n=m$, when each column of the matrix ${\bf D}$ belongs to one of $r$ groups of the collinear columns. Use ${\bf k}^{\prime}=\{k_{1},k_{2},\ldots,k_{r-1}\}$ to write
@@
$W({\bf s},{\bf D})=\sum_{{\bf k}=0}^{{\bf K}}\delta_{{\bf S}({\bf k}^{\prime}),k_{r}{\bf C}_{r}}H({\bf S}({\bf k}^{\prime}))\cdot\prod_{p=1}^{r}W(k_{p},{\bf u}_{p}),\quad K_{r}=\max(\lfloor s_{i}/C_{r,i}\rfloor).$ (69)